# Intrinsic Diophantine approximation: a solution to Mahler’s problem

### v1

@@ 3.1. Recall on equidistribution
We fix $b$ and $S$ as in the theorem, and denote by $\ell>0$ the length of an interval $I\subset\mathbb{T}^{1}\setminus S$ By considering a larger constant $C$ in Theorem 3.4 if one must, we may assume that $\frac{C}{2\ell}\in\mathbb{N}$.
@@
Let $q\in\mathbb{N}$ and let $H$ be a subgroup of $\mathbb{Z}_{q}^{*}$. For any $\varepsilon>0$ there exists $\eta_{\varepsilon}>0$ such that, provided that $q$ is large enough,
(6) $H\geq q^{\varepsilon}\quad\Longrightarrow\quad\max_{p\in\mathbb{Z}_{q}^{*}}\Big|\sum_{h\in H}\exp\Big(2\pi i\frac{hp}{q}\Big)\Big|\leq q^{-\eta_{\varepsilon}}\#H.$
We start with a lemma.
Lemma 3.6.
Let $\varepsilon>0$. Then for any $n\in\mathbb{N}$ and any $q\in Q_{b,n}$ with $q\wedge b=1$ and $q$ large enough, there exists $1\leq t\leq\frac{2C}{\ell}$ such that
$o_{q/(q\wedge t)}(b)<\Big(\frac{q}{q\wedge t}\Big)^{\varepsilon}.$
@@
Since $q\in Q_{b,n}$, there exists $p\in\mathbb{Z}_{q}^{*}$ such that $\frac{p}{q}\in S$. As $S$ is $\times b$-invariant, this implies that
@@
Applying the Erdős–Turán inequality (Theorem 3.4) with $A=\frac{2C}{\ell}$, we obtain
@@
For $1\leq h\leq\frac{2C}{\ell}$, write $q^{\prime}=\frac{q}{h\wedge q}$ and $h^{\prime}=\frac{h}{h\wedge q}$. Then
$\Big|\sum_{0\leq k\leq o_{q}(b)-1}\exp\Big(2\pi i\frac{hpb^{k}}{q}\Big)\Big|=\Big|\frac{o_{q}(b)}{o_{q^{\prime}}(b)}\sum_{0\leq k\leq o_{q^{\prime}}(b)-1}\exp\Big(2\pi i\frac{b^{k}h^{\prime}}{q^{\prime}}\Big)\Big|.$
@@
$\frac{1}{o_{q}(b)}\Big|\sum_{0\leq k\leq o_{q}(b)-1}\exp\Big(2\pi i\frac{hpb^{k}}{q}\Big)\Big|=\frac{1}{o_{q^{\prime}}(b)}\Big|\sum_{0\leq k\leq o_{q^{\prime}}(b)-1}\exp\Big(2\pi i\frac{b^{k}h^{\prime}}{q^{\prime}}\Big)\Big|.$
Since $o_{q^{\prime}}(b)\geq{q^{\prime}}^{\varepsilon}$, by Theorem 3.5,
$\frac{1}{o_{q^{\prime}}(b)}\Big|\sum_{0\leq k\leq o_{q^{\prime}}(b)-1}\exp\Big(2\pi i\frac{b^{k}h^{\prime}}{q^{\prime}}\Big)\Big|\leq{q^{\prime}}^{-\eta_{\varepsilon}}.$
@@
$D(U)\leq\frac{\ell}{2}+\frac{2C}{\ell}\,q^{-\eta_{\varepsilon}}.$
@@
Proposition 3.7.
For all sufficiently large $n\in\mathbb{N}$ and every $\varepsilon>0$,
$N_{\varepsilon,n,b}=\#\Big\{b^{n-1}<q\leq b^{n},\ b\wedge q=1:\ \exists\,1\leq t\leq\frac{2C}{\ell}\text{ with }o_{q/(q\wedge t)}(b)\leq q^{\varepsilon}\Big\}\leq b^{2\varepsilon n}.$
This follows from the next lemma.
Lemma 3.8.
For all sufficiently large $n\in\mathbb{N}$ and every $\varepsilon>0$,
$\Big\{b^{n-1}<q\leq b^{n},\ b\wedge q=1:\ o_{q}(b)\leq q^{\varepsilon}\Big\}\leq b^{2\varepsilon n}.$
Proof.
Observe that
$\prod_{k\leq b^{\varepsilon n}}(b^{k}-1)\leq b^{b^{2\varepsilon n}}.$
If $q\leq b^{n}$ satisfies $o_{q}(b)=k\leq b^{\varepsilon n}$, then $q$ divides $b^{k}-1$. Therefore,
$b^{b^{2\varepsilon n}}\geq\prod_{\begin{subarray}{c}q\leq b^{n}\\ b\wedge q=1\\ o_{q}(b)\leq b^{\varepsilon n}\end{subarray}}q\geq N_{\varepsilon,n,b}!.$
Since for large $N$,
$N!\geq\Big(\frac{N}{e}\Big)^{N}=b^{\,N\frac{\log(N/e)}{\log b}},$
it follows, for $n$ large enough, that $N_{\varepsilon,n,b}\leq b^{2\varepsilon n}$. ∎
Write $\widetilde{Q}_{b,n}=\left\{q\in Q_{b,n}:\ q\wedge b=1\right\}.$
Combining Lemma 3.6 and Proposition 3.7, we conclude that for any $\varepsilon>0$ and all sufficiently large $n$,
$\widetilde{Q}_{b,n}\leq b^{2\varepsilon n},$
and therefore
$\lim_{n\to+\infty}\frac{\log\#\widetilde{Q}_{b,n}}{n}=0.$
Fix $q\in Q_{b,n}$ and write $q=b^{k}q^{\prime}$ where $q^{\prime}\wedge q=1.$ Fix $p\in\mathbb{N},$ $p\wedge q=1$ such that $\frac{p}{q}\in\widetilde{K}_{T}.$ Then, as $K_{T}$ is $\times b$ invariant, $\frac{b^{k}p}{q}=\frac{p}{q^{\prime}}\in\widetilde{K}_{T}$ and thus $q^{\prime}\in\widetilde{Q}_{b,n-k}.$ We conclude that, providing that $n$ is chosen large enough,
$\#Q_{b,n}\leq\sum_{0\leq k\leq n}\#\widetilde{Q}_{b,k}\leq b^{3\varepsilon n}.$
This proves Theorem 3.1.
@@
$\dim_{H}\Big\{y\in K:\$$\big\|y-\pi\big((i_{1},\dots,i_{n}),(i_{k+1},\dots,i_{n})_{\infty}\big)\big\|_{\infty}$
@@
$E_{\psi,T,\text{int}}$ $\subset\limsup_{\underline{i}=(i_{1},\dots,i_{n})\in\Lambda^{*}}\bigcup_{0\leq k\leq n-1}B\Big(\pi\big(i_{1},\dots,i_{n},(i_{k+1},\dots,i_{n})_{\infty}\big),$
@@
$\dim_{H}E_{\psi,T,\text{int}}=$
$\dim_{H}\limsup_{\underline{i}=(i_{1},\dots,i_{n})\in\Lambda^{*},\,0\leq k\leq n-1}B\Big(\pi\big(i_{1},\dots,i_{n},(i_{k+1},\dots,i_{n})_{\infty}\big),\eta\Big(\frac{1}{\prod_{1\leq j\leq k}q_{i_{j}}\,(\prod_{k+1\leq j\leq n}q_{i_{j}}-1)}\Big)\Big).$
@@
$\widetilde{\eta}(f_{\underline{i}}(K))\leq\eta\Big(\frac{1}{\prod_{1\leq j\leq k}q_{i_{j}}(\prod_{k+1\leq j\leq n}q_{i_{j}}-1)}\Big)\leq\widehat{\eta}(f_{\underline{i}}(K)).$
@@ 4.1.1. Proof of the upper bound
$1\leq\#\text{\emph{Uper}}(\underline{i})\leq|f_{\underline{i}}(K)|^{-\varepsilon}.$
@@
Set $\varepsilon_{0}=\varepsilon/5$. First observe that for any $\underline{j}\in cl(\underline{i})$ and any $1\leq k\leq n$,
$\pi\big((j_{k+1},\ldots,j_{n})_{\infty}\big)=f_{\underline{i}}\Big(\pi\big((j_{k+1},\ldots,j_{n})_{\infty}\big)\Big).$
@@
Let
$n_{\underline{i}}=\Big\lfloor\frac{-\log c_{\underline{i}}}{\log 2}\Big\rfloor+1,\qquad 0\leq p\leq n_{\underline{i}}.$
@@
$\widetilde{\text{\emph{Per}}}(\underline{i},p)=\Big\{\pi((j_{k+1},\ldots,j_{n})_{\infty}):(j_{1},\ldots,j_{n})\in cl(\underline{i}),\ 1\leq k\leq n,\ 2^{-(p+1)}<c_{(j_{k+1},\ldots,j_{n})}\leq 2^{-p}\Big\}.$
Note that if $f_{(\widetilde{j}_{k+1},\ldots,\widetilde{j}_{n})}=f_{(j_{k+1},\ldots,j_{n})}$, then
$\pi((\widetilde{j}_{k+1},\ldots,\widetilde{j}_{n})_{\infty})=\pi((j_{k+1},\ldots,j_{n})_{\infty}).$
Therefore,
$\widetilde{\text{\emph{Per}}}(\underline{i},p)\leq\#\Big\{f_{(j_{k+1},\ldots,j_{n})}:\underline{j}\in cl(\underline{i}),\ 1\leq k\leq n,\ 2^{-(p+1)}<c_{(j_{k+1},\ldots,j_{n})}\leq 2^{-p}\Big\}.$
Since for any $\underline{j}\in cl(\underline{i})$,
$2^{-n_{\underline{i}}}\leq c_{\underline{i}}=c_{\underline{j}}\leq 2^{-(n_{\underline{i}}-1)},$
we obtain
$2^{-(n_{\underline{i}}-p)}\leq c_{(j_{k+1},\ldots,j_{n})}\leq 2^{-(n_{\underline{i}}-p-1)}.$
Moreover, if $f_{\underline{j}}(K)=f_{\underline{i}}(K)$, then
@@
By the AWSC, there exists $N_{\varepsilon_{0}}$ such that for all $n\geq N_{\varepsilon_{0}}$,
(8) $\Big\{f_{(j_{1},\ldots,j_{k})}:\underline{j}\in cl(\underline{i}),\ 1\leq k\leq n\Big\}\leq c_{\underline{i}}^{-\varepsilon_{0}}.$
Let
$\mathcal{J}_{p}=\Big\{\widetilde{j}\in cl(\underline{i}):\exists k_{\widetilde{j}}\ \text{s.t. }\ \pi((\widetilde{j}_{k_{\widetilde{j}}+1},\ldots,\widetilde{j}_{n})_{\infty})\in\widetilde{\text{\emph{Per}}}(\underline{i},p)\Big\}.$
If $f_{(j_{1},\ldots,j_{n})}(K)=f_{\underline{i}}(K)$, then
$f_{(j_{k+1},\ldots,j_{n})}(K)=f_{(j_{1},\ldots,j_{k})}^{-1}(f_{\underline{i}}(K)),$
and hence
$f_{(j_{k+1},\ldots,j_{n})}(K)\cap\bigcup_{\widetilde{j}\in\mathcal{J}_{p}}f_{(\widetilde{j}_{1},\ldots,\widetilde{j}_{k_{\widetilde{j}}})}^{-1}(f_{\underline{i}}(K))\neq\varnothing.$
For any $\widetilde{j}\in\mathcal{J}_{p}$,
$\gamma^{-1}2^{-p}\leq\Big|f_{(\widetilde{j}_{1},\ldots,\widetilde{j}_{k_{\widetilde{j}}})}^{-1}(f_{\underline{i}}(K))\Big|\leq\gamma 2^{-p},$
for some constant $\gamma\geq 1$. By (8), the above union can be covered by at most $c_{\underline{i}}^{-2\varepsilon_{0}}$ balls of radius $2^{-p}$, say $L_{1},\ldots,L_{N}$.
Since $2^{-(p-1)}<c_{(j_{k+1},\ldots,j_{n})}\leq 2^{-p}$, Definition 2.4 yields, for each $1\leq\ell\leq N$,
$\Big\{f_{(j_{k+1},\ldots,j_{n})}:\pi((j_{k+1},\ldots,j_{n})_{\infty})\in\widetilde{\text{\emph{Per}}}(\underline{i},p),\ f_{(j_{k+1},\ldots,j_{n})}(K)\cap L_{\ell}\neq\varnothing\Big\}\leq c_{\underline{i}}^{-\varepsilon_{0}}.$
Consequently,
$\widetilde{\text{\emph{Per}}}(\underline{i},p)\leq c_{\underline{i}}^{-3\varepsilon_{0}}.$
Summing over $0\leq p\leq n_{\underline{i}}$ gives
$\text{\emph{Uper}}(\underline{i})\leq n_{\underline{i}}c_{\underline{i}}^{-3\varepsilon_{0}}\leq c_{\underline{i}}^{-4\varepsilon_{0}}.$
Since $|f_{\underline{i}}(K)|=c_{\underline{i}}|K|$, for $n$ large enough,
$\text{\emph{Uper}}(\underline{i})\leq|f_{\underline{i}}(K)|^{-5\varepsilon_{0}}=|f_{\underline{i}}(K)|^{-\varepsilon}.$
@@
$\Lambda_{B}=\Big\{\underline{i}=(i_{1},\ldots,i_{n})\in\Lambda^{*}:c_{\underline{i}}\leq r<c_{(i_{1},\ldots,i_{n-1})}\Big\}.$
@@
    for every $n\geq N_{\varepsilon}$ there exists a covering $\mathcal{C}_{n}$ of $K$ by balls of radius $\lambda^{n}$ satisfying
@@
    for every ball $B$ with $|B|\leq\lambda^{N_{\varepsilon}}$,
@@ 4.1.2. Proof of the lower bound
We recall the following result, established as [10, Proposition 6.7].

### v3

@@ 3.3. Recall on equidistribution
We fix $b$ and $S$ as in the theorem, and denote by $\ell>0$ the length of an interval $I\subset\mathbb{T}^{1}\setminus S$. By considering a larger constant $C$ in Theorem 3.7 if necessary, we may assume that $\frac{C}{2\ell}\in\mathbb{N}$.
@@
Let $q\in\mathbb{N}$ be an integer and let $H$ be a subgroup of $\mathbb{Z}_{q}^{*}$. There exists $C>0$ such that, provided that $q$ is large enough,
(9) $H\geq q^{\frac{C}{\sqrt{\log\log\log q}}}\quad\Longrightarrow\quad\max_{p\in\mathbb{Z}_{q}^{*}}\Big|\sum_{h\in H}\exp\Big(2\pi i\frac{hp}{q}\Big)\Big|\leq o_{q\to+\infty}(1)\#H.$
We prove the following proposition.
Proposition 3.9.
Let $\varepsilon>0$. Then for any large enough $n\in\mathbb{N}$ and any $q\in Q_{b,n}(S)$ with $q\wedge b=1$, there exists $1\leq t\leq\frac{2C}{\ell}$ such that
$O_{q/(q\wedge t)}(b)<\Big(\frac{q}{q\wedge t}\Big)^{f(\frac{q}{q\wedge t})},$
where $f(p)=\frac{C}{\sqrt{\log\log\log p}},$ with $C$ as in Theorem 3.8.
Notice that for every $m\in\mathbb{N}$, for every $q$ large enough, $f(\frac{q}{m})\leq 2f(q).$ Moreover, as, for any $q\in\mathbb{N},$ $q\wedge b=1,$ for any $p$ dividing $q$
$O_{q}(b)\leq O_{q/p}(b)\times O_{p}(b)\leq pO_{q/p}(b),$
Proposition 3.9 implies, with the notation involved, that for $q$ large enough,
$O_{q}(b)\leq\frac{\frac{2C}{\ell}}{(q\wedge t)^{f(\frac{q}{q\wedge t})}}q^{f(\frac{q}{q\wedge t})}\leq q^{3f(q)},$
yielding Theorem 3.1.
@@
Since $q\in Q_{b,n}(S)$, there exists $p\in\mathbb{Z}_{q}^{*}$ such that $\frac{p}{q}\in S$. As $S$ is $\times b$-invariant, this implies that
@@
Applying Erdős–Turán ’s inequality (Theorem 3.7) with $A=\frac{2C}{\ell}$, we obtain
@@
For $1\leq h\leq\frac{2C}{\ell}$, set $q^{\prime}=\frac{q}{h\wedge q}$ and $h^{\prime}=\frac{h}{h\wedge q}$. Then
$\Big|\sum_{0\leq k\leq o_{q}(b)-1}\exp\Big(2\pi i\frac{hpb^{k}}{q}\Big)\Big|=\Big|\frac{o_{q}(b)}{o_{q^{\prime}}(b)}\sum_{0\leq k\leq o_{q^{\prime}}(b)-1}\exp\Big(2\pi i\frac{b^{k}ph^{\prime}}{q^{\prime}}\Big)\Big|.$
@@
$\frac{1}{o_{q}(b)}\Big|\sum_{0\leq k\leq o_{q}(b)-1}\exp\Big(2\pi i\frac{hpb^{k}}{q}\Big)\Big|=\frac{1}{o_{q^{\prime}}(b)}\Big|\sum_{0\leq k\leq o_{q^{\prime}}(b)-1}\exp\Big(2\pi i\frac{b^{k}ph^{\prime}}{q^{\prime}}\Big)\Big|.$
Since $o_{q^{\prime}}(b)\geq{q^{\prime}}^{\varepsilon}$, by Theorem 3.8, provided that $q$ is large enough,
$\frac{1}{o_{q^{\prime}}(b)}\Big|\sum_{0\leq k\leq o_{q^{\prime}}(b)-1}\exp\Big(2\pi i\frac{b^{k}ph^{\prime}}{q^{\prime}}\Big)\Big|\leq{q^{\prime}}^{-\eta_{\varepsilon}}.$
@@
$D(U)\leq\frac{\ell}{2}+\Big(\frac{2C}{\ell}\Big)^{1+\eta_{\varepsilon}}\,q^{-\eta_{\varepsilon}}.$
@@ 4.1. Approximation by eventually periodic sequences
$\dim_{H}\Big\{y\in K:\$$\big\|y-\pi\Big((i_{1},\dots,i_{n},(i_{k+1},\dots,i_{n})_{\infty})\Big)\big\|_{\infty}$
@@
$E_{\psi,T,\text{int}}$ $\subset\widetilde{E}_{\eta,T}:=\limsup_{\underline{i}=(i_{1},\dots,i_{n})\in\Lambda^{*}}\bigcup_{0\leq k\leq n-1}B\Big(\pi\big(i_{1},\dots,i_{n},(i_{k+1},\dots,i_{n})_{\infty}\big),$
@@
$\dim_{H}E_{\psi,T,\text{int}}=\dim_{H}\widetilde{E}_{\eta,T}.$
@@
$\widetilde{\eta}(|f_{\underline{i}}(K)|)\leq\eta\Big(\frac{1}{\prod_{1\leq j\leq k}q_{i_{j}}(\prod_{k+1\leq j\leq n}q_{i_{j}}-1)}\Big)\leq\widehat{\eta}(|f_{\underline{i}}(K)|).$
@@ 4.2.1. Proof of the upper bound
$1\leq\#\mathrm{Uper}(\underline{i})\leq|f_{\underline{i}}(K)|^{-\varepsilon}.$
@@
Set $\varepsilon_{0}=\varepsilon/2$. First observe that for any $\underline{j}\in cl(\underline{i})$ and any $1\leq k\leq n$,
$\pi\big((j_{1},...,j_{n}),(j_{k+1},\ldots,j_{n})_{\infty}\big)$ $=\lim_{t\to+\infty}f_{(j_{1},...,j_{n})}\circ f_{(i_{k+1},...,i_{n},...,i_{k+1},...,i_{n})}(0)$
$=f_{\underline{i}}\Big(\pi\big((j_{k+1},\ldots,j_{n})_{\infty}\big)\Big).$
@@
Moreover, if $(\widetilde{j}_{1},...,\widetilde{j}_{n}),(j_{1},..,j_{n})\in\mathcal{C}l(\underline{i})$ are such that
$f_{(\widetilde{j}_{1},....,\widetilde{j}_{k})}=f_{(j_{1},...,j_{k})},$
one has $f_{(\widetilde{j}_{k+1},....,\widetilde{j}_{n})}=f_{(j_{k+1},...,j_{n})},$ so that
$\pi\big((\widetilde{j}_{k+1},\ldots,\widetilde{j}_{n})_{\infty}\big)=\pi\big((j_{k+1},\ldots,j_{n})_{\infty}\big).$
This yields
$\text{\emph{Uper}}(\underline{i})\leq\#\Big\{f_{(j_{1},...,j_{k})}:\underline{j}=(j_{1},...,j_{n})\in cl(\underline{i}),\ 1\leq k\leq n\Big\}.$
Fix $0<t<c_{1}$ and set
$n_{\underline{i}}=\Big\lfloor\frac{-\log c_{\underline{i}}}{\log t}\Big\rfloor+1,\qquad 0\leq p\leq n_{\underline{i}}.$
@@
$\widetilde{\text{\emph{Per}}}(\underline{i},p)=\Big\{f_{(j_{1},...,j_{k})}:(j_{1},\ldots,j_{n})\in cl(\underline{i}),\ 1\leq k\leq n,\ t^{-(p+1)}<c_{(j_{1},\ldots,j_{k})}\leq t^{-p}\Big\}.$
Fix $x\in f_{\underline{i}}(K)$ and remark that for any $\underline{j}\in\mathcal{C}l(\underline{i}),$
@@
As
$\widetilde{\text{\emph{Per}}}(\underline{i},p)\subset\big\{f_{\underline{i}}:\ \underline{i}\in\Lambda_{|K|t^{p}}\text{ and }f_{\underline{i}}(K)\cap B(x,|K|t^{p})\neq\emptyset\big\},$
by Definition 2.4, there exists a constant $C_{\varepsilon_{0}}>0$ such that
$\widetilde{\text{\emph{Per}}}(\underline{i},p)\leq C_{\varepsilon_{0}}t^{-\varepsilon_{0}p}\leq\frac{C_{\varepsilon_{0}}}{|K|}|f_{\underline{i}}(K)|^{-\varepsilon_{0}}.$
Since
$\bigcup_{0\leq p\leq n_{\underline{i}}}\widetilde{\text{\emph{Per}}}(\underline{i},p)=\widetilde{\text{\emph{Per}}}(\underline{i}),$
one has
$\widetilde{\text{\emph{Per}}}(\underline{i},p)\leq(n_{\underline{i}}+1)\frac{C_{\varepsilon_{0}}}{|K|}|f_{\underline{i}}(K)|^{-\varepsilon_{0}}\leq|f_{\underline{i}}(K)|^{-2\varepsilon_{0}}$
provided that $n$ was chosen large enough. ∎
@@
$\Lambda_{B}=\Big\{\underline{i}=(i_{1},\ldots,i_{n})\in\Lambda^{*}:f_{\underline{i}}(K)\cap B\neq\emptyset,c_{\underline{i}}\leq r<c_{(i_{1},\ldots,i_{n-1})}\Big\}.$
@@
    for every $n\geq N_{\varepsilon}$ (recalling (7)) there exists a covering $\mathcal{C}_{n}$ of $K$ by balls of radius $\lambda^{n}$ satisfying
@@
    for every ball $B$ with $|B|\leq\lambda^{N_{\varepsilon}}$ (recalling that $T$ satisfies AWSC (Definition 2.4),
@@ 4.2.2. Proof of the lower bound
We recall the following result, established as [11, Proposition 6.7].
@@ 5. Some perspectives
5.1. About the connection between equidistributuion mod q and Mahler’s problem
A careful reader will notice, that in order to prove Theorem 1.5, one would only need to prove Conjecture 1.4 for mappings $f_{\epsilon}(q)=\frac{\varepsilon}{\sqrt{\log\log\log q}},$ where $\varepsilon>0,$ this rate being given Theorem 3.8. Thus a more precise conjecture, yielding the conclusion of Theorem 1.5, is the following.
Conjecture 5.1.
Let $b\geq 2$ be an integer. There exists a non increasing mapping $f:\mathbb{N}\to\mathbb{R}_{+}$ such that:
  • (1)
    for every $q\in\mathbb{N},$ $q\wedge b=1,$
    $\Big(O_{q}(b)\geq q^{f(q)}\Big)\Rightarrow\Big(\Big|\max_{p\in\mathbb{Z}_{q}^{*}}\sum_{0\leq k\leq O_{q}(b)-1}\exp(2i\pi\frac{pb^{k}}{q})\Big|\leq O_{q}(b)o_{q\to+\infty}(1)\Big),$
  • (2)
    one has
    $\lim_{x\to+\infty}\frac{\log\#\left\{1\leq q\leq x,q\wedge b=1:\ O_{q}(b)\leq q^{f(q)}\right\}}{\log x}=0.$